OpenStudy (anonymous):

suppose f[x]=x^3. Find the graph of f [x+5]

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (anonymous):

\[f(x)=x^3\\ f(x+5)=(x+5)^3\]

OpenStudy (solomonzelman):

\(\huge\color{ royalblue }{\large {\bbox[5pt, lightyellow ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline \rm{ Shifts ~~~from~~~ f(x)~~~ to~~~g(x)}&~\rm{c~~~units~~~~} \\ \hline \\f(x)= x^3 ~~~~~\rm{\Rightarrow}~~~~ g(x)= (x \normalsize\color{red}{ -~\rm{c} })^3 &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= x^3 ~~~~~\rm{\Rightarrow}~~~~ g(x)= (x \normalsize\color{red}{ +~\rm{c} })^3&~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= x^3 ~~~~~\rm{\Rightarrow}~~~~ g(x)= x^3 \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= x^3 ~~~~~\rm{\Rightarrow}~~~~ g(x)= x^3 \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

so your function is the shift to the left, by 5 units.